The lengths of the sides of a triangle are consecutive odd integers. Find the length of the shortest side if the perimeter is 19 units shorter than four times the length of that side
Answer by elima(1433)

The lengths of the sides of a triangle are consecutive odd integers. Find the length of the shortest side if the perimeter is 19 units shorter than four times the length of that side
The sides will be;
x
x+2
x+4
The perimeter=4x-19
======================
Add all sides together to get the perimeter;
x+x+2+x+4=4x-19
3x+6=4x-19
3x-4x=-19-6
-x=-25
x=25
So the shortest side = 25
next side = 25+2=27
third side = 25+4=29
:)